# Motion-Aware Transformer for Multi-Object Tracking

### v1

@@ 4.1 Dataset and Metrics
We evaluate our approach on three challenging benchmarks: DanceTrack Sun et al. (2022), SportsMOT Cui et al. (2023), and BDD100k Yu et al. (2020).
@@
DanceTrack is a large-scale dataset specifically designed for multi-human tracking in dance scenarios. It is characterized by uniform appearances and diverse, complex motions, which make object association across frames highly challenging. The dataset contains 100 videos, divided into 40 training, 25 validation, and 35 test sequences, with an average video length of 52.9 seconds. Due to its challenging poses and association difficulty, DanceTrack has become one of the key benchmarks for evaluating end-to-end MOT methods.
@@
SportsMOT is a large-scale dataset densely annotated with every player on the court. It contains 240 videos with more than 150K frames across three sports categories, split into 45 training, 45 validation, and 150 test sequences. The dataset includes a wide variety of dynamic scenes with both moving players and camera motion. A key characteristic is that the test set is substantially larger than the training set, which we leverage to complement the limitations of DanceTrack in terms of scale and diversity.
@@
BDD100k is a large-scale driving video dataset consisting of 100K videos with 10 object categories. Its tracking subset includes 1,400 training, 200 validation, and 400 test videos covering 8 valid categories. BDD100k is the largest and most challenging MOT dataset to date, with highly diverse scenarios in terms of object scales, weather conditions, and times of day. We employ this dataset to comprehensively evaluate the generalization ability of our method, particularly in multi-category settings.
@@
For evaluation, we primarily use the Higher Order Tracking Accuracy (HOTA) metric Luiten et al. (2020), which provides a balanced assessment of tracking by decomposing results into detection accuracy (DetA) and association accuracy (AssA). We additionally report the widely used MOTA and IDF1 metrics to give a broader view of detection and identification performance. For multi-category scenarios, we adopt Track Every Thing Accuracy (TETA) Li et al. (2022), which includes sub-metrics for localization accuracy (LocA) and association accuracy (AssocA). TETA has been shown to provide a more reliable and comprehensive evaluation than traditional metrics, especially on large-scale multi-class datasets such as BDD100k.
Table 1: Ablation studies on MATR under different components and settings.
(a) Effect of MATR components. Bl-IMP denotes baseline improvement; MAT denotes inclusion of the MAT module.
@@
(b) Effect of the number of Decoder layers in MAT. ”Dec” indicates the number of Decoder layers. ”KLF” refers to Kalman Filter.
@@
4.2 Implementation Details
We build MATR upon MOTR Zeng et al. (2022) and employ several data augmentation strategies, including random resizing and random cropping. The shorter side of each input image is resized to 800 pixels, with the longer side restricted to a maximum of 1536 pixels. Unless otherwise specified, we use ResNet50 as the backbone network. Additionally, we evaluate a larger backbone, Swin-Tiny Liu et al. (2021), to provide a more comprehensive comparison.
For hyperparameters, the number of detection queries $N_{\det}$ is set to 300, the sequence length $S$ is fixed at 5, and the dropout probability for track queries is 0.1. The batch size is 1 per GPU, with each batch containing a video clip of $S$ frames sampled at random intervals from 1 to 10 frames. We adopt the AdamW optimizer, using an initial learning rate of $2\times 10^{-4}$ for ResNet50 and $1\times 10^{-4}$ for Swin-Tiny. During training, tracked targets with IoU scores below the threshold $\tau_{\text{IoU}}=0.5$ are filtered out. These training settings remain consistent across all experiments.
For the CrowdHuman dataset, we follow CenterTrack Zhou et al. (2020) and apply random spatial shifts to generate pseudo trajectories when including it in training. The overall MATR loss is defined as:
$\mathcal{L}_{\text{MATR}}=\tau_{\text{traj}}\mathcal{L}_{\text{traj}}+\mathcal{L}_{\text{MOTR}},$ (3)
where $\tau_{\text{traj}}=5$ is a fixed weighting coefficient, and $\mathcal{L}_{\text{MOTR}}$ is retained exactly as in MOTR. Our strengthened baseline results are reported in Table 1(a) and serve as the foundation for our experiments. All ablation studies are conducted on the DanceTrack dataset. Final results are reported on the test sets of DanceTrack and SportsMOT, and on the validation set of BDD100k.
Training is performed on 8 NVIDIA A5000 GPUs (24GB each). The ablation studies span 5 epochs, requiring approximately 1 day of training. Detailed ablation results are discussed in Section 4.3. The final MATR model uses the MAT module with one Decoder layer. MATR introduces negligible computational overhead compared to MOTR: +1M parameters and +5% FLOPs, while improving HOTA by more than 9 points. In contrast, MOTRv2/v3 require ¿2x parameters and higher runtime cost, highlighting the efficiency of our design.
Following standard practice in MOTR, CO-MOT, and MeMOTR Zhang et al. (2023); Yu et al. (2023); Yan et al. (2023), we combine DanceTrack training data with CrowdHuman. Training lasts 20 epochs with a learning rate decay of 0.1 applied every 8 epochs, requiring about 2.5 days in total. For SportsMOT, training follows the same setup as DanceTrack but excludes additional datasets. For BDD100k, the reduced input size and classification head modification match the settings used in prior end-to-end baselines, ensuring comparability. We also follow the SportsMOT setup, except training runs for 13 epochs with a learning rate decay at epoch 11. Training on BDD100k takes approximately 7 days.
4.3 Ablation Study
MAT is a novel module designed to replace the QIM Zeng et al. (2022) module used in MOTR and MOTRv2. Unlike QIM, MAT leverages additional input from the Transformer Encoder and applies direct supervision to the predicted trajectories. To minimize the risk of overfitting, all ablation studies are conducted on the DanceTrack dataset, as it provides extensive training data and highly challenging association scenarios that clearly demonstrate the effectiveness of our method.
The ablation results are summarized in Table 1(a). Baseline improvements alone already increase HOTA by 4.6 points compared to MOTR. When replacing QIM with the proposed MAT module (with one Decoder layer), we achieve consistent and substantial improvements across all evaluation metrics. Specifically, compared to MOTR, HOTA improves by 9.4 points and MOTA by 6.7 points. Detection accuracy (DetA) increases modestly by 3.6 points, while the association metrics (AssA and IDF1) improve significantly, by 12.5 and 14.9 points, respectively. Because MOT performance is particularly sensitive to association quality, these results underscore the importance of MAT in enhancing tracking stability.
@@ 4.3 Ablation Study
Since MAT predicts future positions in advance, it shares some conceptual similarity with the Kalman Filter (Welch & Bishop, 1995) (KLF), a classic prediction mechanism used in traditional tracking-by-detection methods. We therefore include KLF as a baseline in Table 1(b). The results indicate that while association accuracy (AssA) increases whenever a prediction mechanism is applied, detection accuracy (DetA) drops sharply when using KLF, indicating that its poor prediction quality prevents high-quality detections from being matched. This shows that KLF’s prediction accuracy is insufficient for end-to-end MOT. The same limitation has long been observed in tracking-by-detection pipelines: their methods typically achieve higher detection accuracy, but the low-quality predictions from KLF often prevent high-quality detections from being matched during association, ultimately degrading performance. In end-to-end methods, this issue is further magnified - poor predictions from KLF directly propagate into the Transformer Decoder, causing a substantial decline in detection accuracy.
@@
Table 2: Comparison with state-of-the-art methods on the DanceTrack test set. ”E2E” indicates whether the method is fully end-to-end. ”Params” denotes the total number of parameters in millions (M). Higher values are better for all metrics. $\ddagger$} indicates results obtained by additionally training on the DanceTrack validation set. The MATR configuration corresponds to the final model described in Section 4.2.
Methods Attributes Metrics
E2E Params(M) HOTA DetA AssA MOTA IDF1
FairMOT Zhang et al. (2021) / 39.7 66.7 23.8 82.2 40.8
CenterTrack Zhou et al. (2020) / 41.8 78.1 22.6 86.8 35.7
TraDeS Wu et al. (2021) / 43.3 74.5 25.4 86.2 41.2
ByteTrack Zhang et al. (2022) / 47.7 71.0 32.1 89.6 53.9
GTR Zhou et al. (2022) / 48.0 72.5 31.9 84.7 50.3
QDTrack Pang et al. (2021) / 54.2 80.1 36.8 87.7 50.4
OC-SORT Cao et al. (2023) / 55.1 80.3 38.3 92.0 54.6
C-BIoU Yang et al. (2023) / 60.6 81.3 45.4 91.6 61.6
TransTrack (Sun et al., 2020) / 45.5 75.9 27.5 88.4 45.2
MOTR Zeng et al. (2022) 40 54.2 73.5 40.2 79.7 51.5
MOTRv3-R50 (Yu et al., 2023) 40 63.9 76.7 53.5 86.8 67.2
MeMOTR Gao & Wang (2023) 48 68.5 80.5 58.4 89.9 71.2
CO-MOT (Yan et al., 2023) 40 69.4 82.1 58.9 91.2 71.9
MATR-R50(ours) 42 69.4 81.5 59.1 91.0 72.6
MOTRv2 Zhang et al. (2023) 94 69.9 83.0 59.0 91.9 71.7
MOTRv3 (Yu et al., 2023) 103 70.4 83.8 59.3 92.9 72.3
MATR(ours) 43 71.3 82.6 61.6 91.9 75.3
MATR(ours) 43 73.9 84.5 64.8 92.9 76.5
@@ 4.4 Comparison on DanceTrack
We evaluate our final MATR model on the DanceTrack dataset, which contains complex motion scenarios that are particularly challenging for classical linear motion models. MATR delivers significant improvements in both association accuracy and overall tracking performance. Table 2 reports a comparison between MATR and other state-of-the-art methods on the DanceTrack test set. Our method achieves a new state-of-the-art HOTA score of 71.3, with particularly strong association metrics, including 61.6 AssA and 75.3 IDF1.
Due to the inherent limitations of linear motion prediction, tracking-by-detection methods often excel at detection accuracy but fall short in association. For example, OC-SORT Cao et al. (2023) achieves excellent detection accuracy (DetA = 80.3) but struggles with association in complex motion scenarios (AssA = 38.3). In contrast, MATR improves both detection and association metrics simultaneously, highlighting the benefit of explicitly modeling motion to reduce query collisions.
For fairness, we also report MATR results with a ResNet50 backbone, denoted as MATR-R50, aligned with MOTR, MeMOTR Gao & Wang (2023), and CO-MOT Yan et al. (2023). We additionally include MOTRv3-R50 (Yu et al., 2023) for reference. Compared with other ResNet50-based methods, MATR achieves a HOTA score of 69.4, matching CO-MOT. Notably: 1) MATR improves HOTA by 15.2 points over MOTR, together with substantial gains across all other metrics. 2) MATR surpasses MeMOTR by 0.9 HOTA and consistently achieves stronger association performance, without relying on merging strategies across historical frames. 3) MATR matches CO-MOT’s HOTA performance despite not employing explicit detection enhancements, trailing slightly in DetA (-0.6) but outperforming in AssA (+0.2) and IDF1 (+0.7). These findings demonstrate that the consistent training strategies introduced by MAT not only boost detection accuracy but also yield even greater improvements in association quality.
We further compare MATR with larger and more complex models such as MOTRv2 Zhang et al. (2023) and MOTRv3 Yu et al. (2023). MATR achieves a HOTA of 71.3, surpassing MOTRv3 by 0.9 points, while maintaining an elegant and concise end-to-end structure. In contrast, MOTRv2 incorporates YOLOX as an external detector, and MOTRv3 employs multiple complex strategies. Despite their complexity, these methods rely on significantly larger backbones (94M and 103M parameters, respectively), whereas MATR achieves superior performance with a smaller Swin-Tiny backbone (43M parameters). Although MATR trails MOTRv3 slightly in DetA (-1.2) and MOTA (-1.0) due to the absence of specialized detection enhancement techniques, it achieves state-of-the-art association results (+2.3 AssA and +3.0 IDF1). This confirms the robustness and effectiveness of explicitly addressing query collisions.
Table 3: Comparison with state-of-the-art methods. Higher values indicate better performance across all metrics.
(a) Results on the SportsMOT test set.
@@
(b) Results on the BDD100k validation set.

### v2

@@ 4.1 Dataset and Metrics
We evaluate our approach on three challenging benchmarks: DanceTrack [sun2022dancetrack], SportsMOT [cui2023sportsmot], and BDD100k [bdd100k].
@@
DanceTrack is a large-scale dataset specifically designed for multi-human tracking in dance scenarios. It is characterized by uniform appearances and diverse, complex motions, which make object association across frames highly challenging. Due to its challenging poses and association difficulty, DanceTrack has become one of the key benchmarks for evaluating end-to-end MOT methods.
@@
SportsMOT is a large-scale dataset densely annotated with every player on the court. The dataset includes a wide variety of dynamic scenes with both moving players and camera motion. A key characteristic is that the test set is substantially larger than the training set, which we leverage to complement the limitations of DanceTrack in terms of scale and diversity.
@@
BDD100k is a large-scale driving video dataset. Its tracking subset includes 1,400 training and 200 validation covering 8 valid categories. BDD100k is the largest and most challenging MOT dataset to date, with highly diverse scenarios in terms of object scales, weather conditions, and times of day. We employ this dataset to comprehensively evaluate the generalization ability of our method, particularly in multi-category settings.
Methods Attributes Metrics
E2E P(M) HOTA DetA AssA MOTA IDF1
FairMOT [zhang2021fairmot] / 39.7 66.7 23.8 82.2 40.8
CenterTrack [zhou2020tracking] / 41.8 78.1 22.6 86.8 35.7
TraDeS [wu2021track] / 43.3 74.5 25.4 86.2 41.2
ByteTrack [zhang2022bytetrack] / 47.7 71.0 32.1 89.6 53.9
GTR [zhou2022global] / 48.0 72.5 31.9 84.7 50.3
QDTrack [pang2021quasi] / 54.2 80.1 36.8 87.7 50.4
OC-SORT [cao2023observation] / 55.1 80.3 38.3 92.0 54.6
C-BIoU [yang2023hard] / 60.6 81.3 45.4 91.6 61.6
TransTrack [sun2020transtrack] / 45.5 75.9 27.5 88.4 45.2
MOTR [zeng2022motr] 40 54.2 73.5 40.2 79.7 51.5
MOTRv3-R50 [yu2023motrv3] 40 63.9 76.7 53.5 86.8 67.2
MeMOTR [gao2023memotr] 48 68.5 80.5 58.4 89.9 71.2
CO-MOT [yan2023bridging] 40 69.4 82.1 58.9 91.2 71.9
MATR-R50(ours) 42 69.4 81.5 59.1 91.0 72.6
MOTRv2 [zhang2023motrv2] 94 69.9 83.0 59.0 91.9 71.7
MOTRv3 [yu2023motrv3] 103 70.4 83.8 59.3 92.9 72.3
MATR(ours) 43 71.3 82.6 61.6 91.9 75.3
MATR(ours) 43 73.9 84.5 64.8 92.9 76.5
Table 1: Comparison with state-of-the-art methods on the DanceTrack test set. ”E2E” indicates whether the method is fully end-to-end. ”P” denotes the total number of parameters in millions (M). Higher values are better for all metrics. $\ddagger$} indicates results obtained by additionally training on the DanceTrack validation set.
@@
For evaluation, we primarily use the Higher Order Tracking Accuracy (HOTA) metric [luiten10higher], which provides a balanced assessment of tracking by decomposing results into detection accuracy (DetA) and association accuracy (AssA). We additionally report the widely used MOTA and IDF1 metrics to give a broader view of detection and identification performance. For multi-category scenarios, we adopt Track Every Thing Accuracy (TETA) [li2022tracking], which includes sub-metrics for localization accuracy (LocA) and association accuracy (AssocA). TETA has been shown to provide a more reliable and comprehensive evaluation than traditional metrics, especially on large-scale multi-class datasets such as BDD100k.
4.2 Implementation Details
We build MATR upon MOTR [zeng2022motr] and employ several data augmentation strategies, including random resizing and random cropping. The shorter side of each input image is resized to 800 pixels, with the longer side restricted to a maximum of 1536 pixels. Following [zeng2022motr], we use ResNet50 as the backbone network. Additionally, we evaluate a larger backbone, Swin-Tiny [liu2021swin], to provide a more comprehensive comparison. For hyperparameters, the detection queries $N_{\det}$ is set to 300, the sequence length $S$ is fixed at 5, and the dropout probability for track queries is 0.1. The batch size is 1 per GPU, with each batch containing a video clip of $S$ frames sampled at random intervals from 1 to 10 frames. We adopt the AdamW optimizer, using an initial learning rate of $2\times 10^{-4}$ for ResNet50 and $1\times 10^{-4}$ for Swin. Tracked targets with IoU scores below the threshold $\tau_{\text{IoU}}=0.5$ are filtered out. These training settings remain consistent across all experiments. For the CrowdHuman dataset, we follow CenterTrack [zhou2020tracking] and apply random spatial shifts to generate pseudo trajectories when including it in training. The overall MATR loss is defined as:
$\mathcal{L}_{\text{MATR}}=\tau_{\text{traj}}\mathcal{L}_{\text{traj}}+\mathcal{L}_{\text{MOTR}}$ (3)
where $\tau_{\text{traj}}=5$ is a fixed weighting coefficient, and $\mathcal{L}_{\text{MOTR}}$ is retained exactly as in MOTR. All ablation studies are conducted on the DanceTrack dataset spanning 5 epochs. Training is performed on 8 NVIDIA A5000 GPUs (24GB each). The final MATR model uses the MAT module with one Decoder layer. MATR introduces negligible computational overhead compared to MOTR: +1M parameters and +5% FLOPs, while improving HOTA by more than 9 points. In contrast, MOTRv2/v3 require $2\times$ parameters and higher runtime cost, highlighting the efficiency of our design.
Following standard practice in MOTR, CO-MOT, and MeMOTR [zhang2023motrv2, yu2023motrv3, yan2023bridging], we combine DanceTrack training data with CrowdHuman. Training lasts 20 epochs with a learning rate decay of 0.1 applied every 8 epochs, requiring about 2.5 days in total. For SportsMOT, training follows the same setup as DanceTrack but excludes additional datasets. For BDD100k, the reduced input size and classification head modification match the settings used in prior end-to-end baselines, ensuring comparability. We also follow the SportsMOT setup, except training runs for 13 epochs with a learning rate decay at epoch 11, takes approximately 7 days.
@@ 4.2 Implementation Details
Table 2: Effect of MATR components. Bl-IMP denotes baseline improvement; MAT denotes inclusion of the MAT module.
4.3 Ablation Study
Unlike QIM, MAT leverages additional input from the Transformer Encoder and applies direct supervision to the predicted trajectories. We conduct ablation studies on the DanceTrack dataset, as it provides extensive training data and highly challenging association scenarios that clearly demonstrate the effectiveness of our method. The ablation results are summarized in Table 2. Baseline improvements alone already increase HOTA by 4.6 points compared to MOTR. When replacing QIM with the proposed MAT module (with one Decoder layer), we achieve consistent and substantial improvements across all evaluation metrics. Specifically, compared to MOTR, HOTA improves by 9.4 points and MOTA by 6.7 points. Detection accuracy (DetA) increases modestly by 3.6 points, while the association metrics (AssA and IDF1) improve significantly, by 12.5 and 14.9 points, respectively. Because MOT performance is particularly sensitive to association quality, these results underscore the importance of MAT in enhancing tracking stability.
@@ 4.3 Ablation Study
Table 3: Effect of the number of Decoder layers in MAT. ”Dec” indicates the number of Decoders. ”KLF” refers to Kalman Filter.
@@
Since MAT predicts future positions in advance, it shares some conceptual similarity with the Kalman Filter [welch1995kalman] (KLF), a classic prediction mechanism used in traditional tracking-by-detection methods. We therefore include KLF as a baseline in Table 3. The results indicate that while association accuracy (AssA) increases whenever a prediction mechanism is applied, detection accuracy (DetA) drops sharply when using KLF, indicating that its poor prediction quality prevents high-quality detections from being matched. This shows that KLF’s prediction accuracy is insufficient for end-to-end MOT. The same limitation has long been observed in tracking-by-detection pipelines: their methods typically achieve higher detection accuracy, but the low-quality predictions from KLF often prevent high-quality detections from being matched during association, ultimately degrading performance. In end-to-end methods, this issue is further magnified - poor predictions from KLF directly propagate into the Transformer Decoder, causing a substantial decline in detection accuracy.
@@ 4.4 Comparison on DanceTrack
MATR delivers significant improvements in both association accuracy and overall tracking performance. Table 1 reports a comparison between MATR and other SOTA methods on the DanceTrack test set. Our method achieves a new SOTA HOTA score of 71.3, with particularly strong association metrics, including 61.6 AssA and 75.3 IDF1. Due to the inherent limitations of linear motion prediction, tracking-by-detection methods often excel at detection accuracy but fall short in association. For example, OC-SORT [cao2023observation] achieves excellent detection accuracy (DetA = 80.3) but struggles with association in complex motion scenarios (AssA = 38.3). MATR improves both detection and association metrics simultaneously, highlighting the benefit of explicitly modeling motion to reduce query collisions.
For fairness, we also report MATR results with a ResNet50 backbone aligned with [zeng2022motr, gao2023memotr, yan2023bridging] , denoted as MATR-R50. We additionally include MOTRv3-R50 [yu2023motrv3] for reference. Compared with other ResNet50-based methods, MATR achieves a HOTA score of 69.4, matching CO-MOT. Notably: (i) MATR improves HOTA by 15.2 points over MOTR, together with substantial gains across all other metrics; (ii) MATR surpasses MeMOTR by 0.9 HOTA and consistently achieves stronger association performance, without relying on merging strategies across historical frames; (iii) MATR matches CO-MOT’s HOTA performance despite not employing explicit detection enhancements, trailing slightly in DetA (-0.6) but outperforming in AssA (+0.2) and IDF1 (+0.7). These findings demonstrate that the consistent training strategies introduced by MAT not only boost detection accuracy but also yield even greater improvements in association quality.
@@
Table 4: Comparison with SOTA methods on the SportsMOT.
We further compare MATR with larger and more complex models such as MOTRv2/3 [zhang2023motrv2, yu2023motrv3]. MATR achieves a HOTA of 71.3, surpassing MOTRv3 by 0.9 points, while maintaining an elegant and concise end-to-end structure. In contrast, MOTRv2 incorporates YOLOX as an external detector, and MOTRv3 employs multiple complex strategies. Despite their complexity, these methods rely on significantly larger backbones (94M and 103M parameters, respectively), whereas MATR achieves superior performance with a smaller Swin-Tiny backbone (43M parameters). Although MATR trails MOTRv3 slightly in DetA (-1.2) and MOTA (-1.0) due to the absence of specialized detection enhancement techniques, it achieves state-of-the-art association results (+2.3 AssA and +3.0 IDF1). This confirms the robustness and effectiveness of explicitly addressing query collisions.
4.5 Comparison on SportsMOT
We evaluate MATR on the SportsMOT [cui2023sportsmot] dataset, Table 4 presents comparisons between MATR and state-of-the-art methods on the SportsMOT test set, all under settings without additional datasets for pretraining or fine-tuning.
Detection performance on SportsMOT is generally strong across methods, as indicated by MOTA scores consistently above 90. However, association accuracy (AssA) remains comparatively limited, typically around or below 60. Since SportsMOT already provides a sufficient amount of training data, we emphasize improving association accuracy rather than relying on external datasets. It is also important to note that SportsMOT annotations cover only athletes and exclude spectators, which differs significantly from datasets such as CrowdHuman [shao2018crowdhuman]. While traditional tracking-by-detection methods can effectively filter irrelevant detections, these annotation inconsistencies introduce additional challenges for end-to-end approaches.
Thanks to the baseline results reported by [gao2023memotr], we are able to perform fair comparisons without external data. MATR significantly improves upon previous methods across all evaluation metrics. In particular, MATR achieves a new state-of-the-art HOTA score of 72.7, representing gains of 2.7 points over MeMOTR [gao2023memotr] and 4.6 points over OC-SORT [cao2023observation]. While our detection accuracy (DetA) improves only slightly compared to OC-SORT (+0.5 points), the most notable gains are observed in association metrics: +2.9 AssA and +3.2 IDF1 compared to MeMOTR. These results highlight the effectiveness of MATR’s motion-aware strategy, which consistently optimizes both tracking and detection queries, leading to substantial improvements in association performance while preserving strong detection accuracy.
4.6 Comparison on BDD100k
As the largest and most diverse MOT dataset, BDD100k [bdd100k] poses substantial challenges for tracking methods. Recently, TETer [li2022tracking] introduced the TETA metric, a comprehensive evaluation framework designed specifically for BDD100k, along with their method. Although TETer achieves the highest localization accuracy (mLocA = 47.2), like other tracking-by-detection methods, it struggles with association metrics, which reduces its overall performance on mTETA and mHOTA.
As shown in Table 5, MATR achieves state-of-the-art performance on nearly all metrics. Compared to MOTR, MATR provides large improvements across the board (+4.0 mTETA, +4.6 mHOTA, +6.0 mLocA, +8.0 mAssocA, and +5.7 mAssA), demonstrating the effectiveness of our MAT module over the original QIM design. For prior SOTA methods such as CO-MOT [yan2023bridging] and MeMOTR [gao2023memotr], which rely on detection-specific tricks or historical frame information, MATR also achieves superior results. Although CO-MOT and MeMOTR both reach similar association scores (56.2 and 56.7 mAssocA, respectively), our method improves by about 3 absolute points, establishing a new state of the art at 59.0 mAssocA.